Calculate the volume of a parallelopiped whose vertices are given by (0,0,0),
(2,1,3),(3,2,4) and (4,2,1).
1
Expert's answer
2020-06-10T07:50:50-0400



(https://en.wikipedia.org/wiki/Triple_product#/media/File:Parallelepiped_volume.svg)



So,


"a=(2,1,3);\nb=(3,2,4);\nc=(4,2,1)."


"V=a\cdot(b\times c)=\begin{vmatrix}
 3 & 2&4 \\
 2 & 1 &3\\
 4&2&1
\end{vmatrix}=5"